Confidence Intervals for Means and Proportions: Problem 9 Previous Problem List Next (1 point) A poll is taken in which 384 out of 525 randomly selected voters indicated their preference for a certain candidate Find a 90% confidence interval for p to Note: You can earn partial credit on this problem. Preview My AnswersSubmit Answers You have attempted this problem 0 times. You have 99 attempts remaining Email instructor
Answer #1

Solution:- Given X = 384 , n = 525

p = X/n = 384/525 = 0.731

q = 1 - p = 1 - p = 1-0.73 = 0.269

90% confidence interval for the Z = 1.645

90% confidence interval for p = p +/- Z*sqrt(pq/n)
= 0.731 +/- 1.645*sqrt(0.731*0.269/525)
= 0.699 , 0.763